OpenStudy (anonymous):

URGENT... What is the slope of a line perpendicular to the line with equation y = 5x – 2? a= –5 b= 5 c= -5/3 (Fraction) d= –2

OpenStudy (anonymous):

NEED HELP DO YOU

OpenStudy (anonymous):

yeah...isn't that why you post on here?

OpenStudy (anonymous):

OK THEN BYE I AM TRYING TO HELP YOU BYE

OpenStudy (anonymous):

I didn't mean it to be rude, my bad.

OpenStudy (anonymous):

NO PROBLEM OK ILL HELP YOU

OpenStudy (anonymous):

Hello, Your equation is in slope-intercept form. \[y = mx +c\] Where \(m\) is your slope and \(c\) is your y-intercept. Do you see how you can apply this knowledge to find your answer?

OpenStudy (anonymous):

Not really.. I'm slightly confused. I'm very, very bad at math :/

OpenStudy (anonymous):

Okay, look at your given equation. \[y = 5x - 2\] Notice how it looks like: \[y = mx + c\] Your question asks for the slope of the given equation. The variable \(m\) is the slope.

OpenStudy (anonymous):

So, 5 would be the slope and 2 would be my y-intercept?

OpenStudy (anonymous):

Close, \(-2\) would be your y-intercept. Other than that you are correct.

OpenStudy (anonymous):

Oh, so it would be opposite? Thank you.

OpenStudy (briensmarandache):

opposite? no mx+b is the standard y intercept you have 5x-2, c=-2

OpenStudy (anonymous):

You are welcome. Not sure what you mean by opposite. The y-intercept is \(-2\) due to it having the negative symbol in front of 2. For example: \[y = 5x + 3\] Slope: 5 Y-intercept: 3 For example: \[y = 2x - 3\] Slope: 2 Y-intercept: -3

OpenStudy (anonymous):

Oh, okay, i think i understand now.

Nnesha (nnesha):

i didn't understand one thing they said perpendicular right ???

Nnesha (nnesha):

perpendicular line should be flip and opposite sign of the equation

Nnesha (nnesha):

if -a is your slope thn positive 1/a suppose to be perpendicular but i didn't see -1/5 in multiple choice because slope is 5 then flipp -1/5 and change sign may be that perpendicular word not suppose to be there or something else